a boat travels 35 km upstream and then came back in a total time of 4 h 48 min. If the speed of the current is 2.5 km/h, what is the speed of the boat?
Answer by checkley79(3341)

D=RT
T=D/R
4.8=35/(R+2.5)+35/(R-2.5)
4.8=[35(R-2.5)+35(R+2.5)]/(R+2.5)(R-2.5)
4.8=[35R-87.5+35R+87.5]/(R^2-6.25)
4.8=70R/(R^2-6.25)
4.8(R^2-6.25)=70R
4.8R^2-30-70R=0
4.8R^2-70R-30=0
(4.8R+2)(R-15)=0
R-15=0
R=15 kmh. ANS FOR THE SPEED OF THE BOAT.
PROOF:
4.8=35/(15+2.5)+35/(15-2.5)
4.8=35/17.5+35/12.5
4.8=2+2.8
4.8=4.8